If a sample of 50 grades from a statistics test has a mean of 81.6 & gives a sample standard deviation of 8.3, what is the 95% comfidence interval for the population mean?
1
Expert's answer
2019-09-17T10:59:10-0400

Confidence Interval for the Mean (CI) = mean (m) ± margin of error (ME)

m = 81.6

"ME=Z*S/\sqrt{n}"

n = 50; S = 8.3; Z95% = 1.96

"ME=1.96*8.3/\sqrt{50}"

"ME=2.30"  

95% CI for the Mean = 81.6 ± 2.30 = (79.3, 83.9)

The 95 percent confidence interval for the mean ranges between 79.3 and 83.9